DIFFERENTIAL GEOMETRY/PDE SEMINAR

Wednesday, June 4, 2003 Padelford C-36 4:00 pm (Note: unusual starting time)

Averages over space curves and associated maximal operators

Malabika Pramanik (University of Wisconsin)

In this work joint with Andreas Seeger, we consider the L^{p} regularity of an averaging operator over a curve in ℝ^{3} with non-vanishing curvature and torsion. We also prove related local smoothing estimates, which lead to L^{p}- boundedness of a certain maximal function associated to these averages. The common thread underlying the proof of these results is a deep theorem of T. Wolff on cone multipliers.

For more information about this seminar, visit the DG/PDE Seminar Web page (from the Math Department home page, www.math.washington.edu, follow the link Seminars, Colloquia, and Conferences).

The University of Washington is committed to providing access, equal opportunity and reasonable accommodation in its services, programs, activities, education and employment for individuals with disabilities. To request disability accommodation contact the Disability Services Office at least ten days in advance at: 206-543-6450/V, 206-543-6452/TTY, 206-685-7264 (FAX), or dso@u.washington.edu.